OpenStudy (anonymous):

Consider the vector v = (0;3;4) in 3D. Find unit vectors u1 and u2 which are perpendicular to v in 3D.

OpenStudy (anonymous):

Do you know the cross product?

OpenStudy (anonymous):

Notice that \(\mathbf{u}_1=(1,0,0)\) is perpendicular. That one is kinda a freebie. Then you wanna do the cross product: \[ \mathbf{v}\times \mathbf{u}_1 = (0,3,4)\times (1,0,0)=\mathbf{w} \]And you make \(\mathbf{w}\) a unit vector: \[ \mathbf{u}_2=\frac{\mathbf{w}}{\|\mathbf{w}\|} \]

OpenStudy (anonymous):

The cross product of two vectors gives you a perpendicular vector.

OpenStudy (anonymous):

I know the cross product, but why u1=(1,0,0) is perpendicular?

OpenStudy (anonymous):

Well you can tell that if you do the dot product it is \(0\).

OpenStudy (anonymous):

\[ \mathbf{v}\cdot \mathbf{u}_1 = (0,3,4)\cdot (1,0,0) = 0\cdot 1+3\cdot 0+4\cdot 0 =0+0+0=0 \]

OpenStudy (anonymous):

Basically if you have some vector like... \[ \mathbf{a} = (a_1, a_2,\dots,a_k,\dots,a_n) \]And you have \(a_k=0\). It will be perpendicular to some vector \[ \mathbf{b} = (b_1, b_2,\dots,b_k,\dots,b_n) \]Where \(b_k=1\) and all other components are \(0\): \(\forall i\ne k:\;b_i=0\)

OpenStudy (anonymous):

Now I am confused as the question says there are u1 and u2.

OpenStudy (anonymous):

Or do we have infinite solutions?

OpenStudy (anonymous):

There are infinite solutions.

OpenStudy (anonymous):



OpenStudy (anonymous):

\(\mathbf{v}\) is unchanged and yet we can rotate our unit vectors an infinitesimal amount.

OpenStudy (anonymous):

Got it. Thank you very much wio!

OpenStudy (anonymous):

Then how to present the answer saying there are infinite solutions in 3D which are perpendicular to v?

OpenStudy (anonymous):

up

OpenStudy (anonymous):

can somebody help me please?